### v1

@@ 5.1 Policy Network Architecture
5.1.1 Input Embedding Module.
Taking the bipartite graph representation of instance $U$ as input, the input embedding module exploits the raw features and topology of the bipartite graph to produce a high-dimensional embedding for each market node and product node. GNN is an effective framework that naturally operates on graph-structured data and learns relational information through message-passing between the nodes on the graph (Zhou et al. 2020, Cappart et al. 2023). Therefore, we leverage GNNs in the input embedding module to embed the bipartite graph and capture the relations between markets and products.
[…]
@@
where $\operatorname{MLP}_{\mathcal{K}}$ and $\operatorname{MLP}_{\mathcal{M}}$ are multi-layer perceptrons (MLPs) for updating the product nodes and market nodes, respectively, and $\mathcal{N}\left(k\right)$ and $\mathcal{N}\left(m\right)$ denote the 1-hop neighbor node set of product node $k$ and market node $m$, respectively. Both phases can be viewed as a modification of the graph isomorphism network (GIN) (Xu et al. 2019) that further takes into account the bipartite structure and incorporates edge embeddings in the update. The updated product node embeddings $g_{k}^{0}$ for $k=1,2,\dots,|K|$ and market node embeddings $h_{m}^{0}$ for $m=0,1,\dots,|M|$ are then used in the market encoder and the decoder.
We remark that our design behind the two-phase embedding procedure is intuitive: in the first phase, we
make each product node collect its supply information, including in which markets it is supplied, the coordinates of these markets, as well as the supply quantities and prices; and in the second phase, we make each market node collect higher-level information about its supplied products, including not only their total demands but also the information about how these products are supplied at other markets, which is extracted based on the product node embeddings from the first phase. Consequently, the product node embedding $g_{k}^{0}$ aggregates its global supply information, and the market node embedding $h_{m}^{0}$ collects more complex information such as the substitutive and complementary relations of product supply with other markets, which is important for the market selection decisions in route construction.
5.1.2 Market Encoder.
It is empirically observed from prior works for routing problems that deeper information extraction for the city nodes, i.e., market nodes in our task, can potentially improve the capability of the policy network (Kwon et al. 2020, Joshi et al. 2022). Thus, we process the market node embeddings through a market encoder for further information extraction.
Following the Transformer (Vaswani et al. 2017) architecture, we process the market node embeddings through $N$ stacked attention layers. Each attention layer consists of two sublayers: an MHA layer that executes information aggregation for the market nodes and a node-wise MLP feed-forward layer to update the aggregated embeddings, where both sublayers add a residual connection (He et al. 2016) and batch normalization (BN) (Ioffe and Szegedy 2015):
@@
The demand context $g_{\mathcal{K}}^{t}$ contains the information of remaining demand $d^{t}$. Third, the route context $o_{\pi}^{t}$ is the embedding for the partial route $\pi^{t}$, i.e., a sequence of market/depot nodes. We adopt an LSTM (Hochreiter and Schmidhuber 1997) recurrent network, similar to the decoding RNN in Ptr-Net (Vinyals et al. 2015), to obtain the route context $o_{\pi}^{t}$:
@@ 5.2 Basic Training Algorithm
The policy network is updated based on the reward received after the solving process. As introduced in Section 3.2.4 and Section 3.2.5, given a TPP instance $U$, the policy network $\theta$ generates the final route $\pi$ with probability $p_{\theta}(\pi\mid U)=\prod_{t=1}^{T}p_{\theta}(a_{t}\mid s_{t})$ during training, where the associated traveling cost can be directly accessed, and the purchasing cost is obtained by solving a transportation problem. To align with the notations in DRL, we let $L(\pi\mid U)$ denote the loss of $\pi$, which is defined as the negative reward for the terminal state. We update the policy network $\theta$ to minimize the expectation of loss $L(\pi\mid U)$, using the REINFORCE algorithm (Williams 1992) with baseline $b(U)$:
@@
where the baseline $b(U)$ can be seen as an estimation for the difficulty of instance $U$ to measure the relative advantage of route $\pi$ generated by the policy network. A well-defined baseline can significantly reduce the variance of gradients and accelerate training. We define the baseline $b(U)$ as the loss of the route obtained from a deterministic greedy rollout, i.e., where the action with the highest probability is always selected at each decision step, based on a baseline network $\theta^{\text{BL}}$ (Kool et al. 2019). Similar to the target Q-network in DQN (Mnih et al. 2015), the baseline network $\theta^{\text{BL}}$ is a copy of policy network $\theta$, but is fixed during each epoch to stabilize the baseline value. The baseline network $\theta^{\text{BL}}$ is periodically replaced by the latest policy network $\theta$ if the improvement of the policy is significant enough according to a paired t-test. Based on the gradient $\nabla_{\theta}\mathcal{L}(\theta\mid U)$, the parameters of policy network are updated using the Adam optimizer (Kingma and Ba 2015). The basic training algorithm introduced above is outlined as Algorithm 1.
@@ 6 Meta-learning Strategy
Unfortunately, the meta-gradient involves a second-order derivative in $\frac{\partial\theta_{\text{in}}^{N}}{\partial\theta_{\text{in}}^{0}}$, which is significantly expensive in computation. Therefore, we instead employ a computationally efficient first-order approximation for the meta-gradient. The update for the meta policy network is approximated as moving towards the parameters of the final inner-loop policy network (Nichol et al. 2018):
@@
Algorithm 2 Meta-training of the policy network
@@ 7.1 Experimental Settings
7.1.1 Instances.
[…]
@@
7.1.2 Baseline Methods.
We select several well-established TPP heuristics reported in (Manerba et al. 2017), including GSH, CAH, and TRH. A widely used and effective practice is to incorporate constructive heuristics (e.g., GSH and CAH) with TRH to remove redundant markets as soon as a solution is produced. Additionally, the solution can potentially be further improved by using a TSP solver to re-sequence the order of visited markets. In our experiment, we employ GSH and CAH, both followed by TRH and the TSP re-sequence, as the baselines, denoted as “GSH+TRH” and “CAH+TRH”, respectively.
Note that we do not include exact methods in the evaluation on synthetic instances since it can take several days to weeks to solve an evaluation instance set using exact methods (even for U-TPP). For a comparison of the optimality gap, we report the best-known solutions in the literature (Riera-Ledesma 2012) on the TPPLIB benchmark instances to assess the optimality gaps of the baseline heuristics and our DRL-based approach.
Table 1: Training configuration.
@@
7.1.3 Training and Inference of the Policy Network.
@@
In the evaluation of our DRL-based approach, we execute the inference of the learned policy network with greedy decoding (similar to the greedy rollout baseline) and instance augmentation (Kwon et al. 2020). The produced route and purchasing plan can be directly used as an end-to-end solution to the TPP instance, denoted as “RL-E2E”. Similar to the baseline heuristics, this end-to-end solution can be further improved through a post-optimization procedure (TRH and TSP re-sequence), denoted as “RL+TRH”. For a fair comparison, the baseline heuristics and our approach are implemented on the same machine and environment.
@@ 7.2 Training Performance
The policy network is trained from scratch for U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$ and (50,100)(50,100)( 50 , 100 ). For brevity, we illustrate the training curve on U-TPP instances of size $\left(|M|,|K|\right)=(50,50)$ in Figure 4(a), which is plotted based on the average loss on the evaluation instance set. The stable convergence demonstrates that our policy network is capable of learning an effective policy for small-sized TPPs when trained from scratch using the basic training algorithm.
However, the policy network suffers serious training collapse when trained from scratch on large-sized U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(100,50)$ and (100,100)(100,100)( 100 , 100 ). As shown in Figure 4(b), the policy network trained from scratch on U-TPP instances of size $\left(|M|,|K|\right)=(100,50)$ is stuck in a high-loss region, and the loss completely fails to decrease throughout training. This is mainly because the state-action space is so large that the policy network cannot find an effective or even reasonable solution if the training starts from a random initialization. Therefore, we instead apply our meta-learning strategy to train the policy network for U-TPP and R-TPP instances with $\left(|M|,|K|\right)=(100,50)$ and (100,100)(100,100)( 100 , 100 ). As shown in Figure 4(b), the utilization of meta-learning strategy significantly promotes the training for an effective policy.
@@ 7.3 Results on Synthetic Instances
Following the training procedures described above, we train the policy network from scratch for U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$ and (50,100)(50,100)( 50 , 100 ), and apply the meta-learning strategy for training U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(100,50)$ and (100,100)(100,100)( 100 , 100 ). Then, the learned policy network is evaluated on the set of 1000 synthetic evaluation instances for each instance distribution. We report the metrics of 1) the average objective value and 2) the runtime (in seconds) for each evaluation instance set. The results are reported in Table 2 for U-TPP and Table 3 for R-TPP.
@@
Table 3: Results on synthetic R-TPP instances.
@@ 7.5 Generalization Performance on Large-Sized Instances
Table 5: Results on larger-sized instances.
@@
Finally, we examine the zero-shot generalization performance of our approach. We use the meta-learning strategy to train the policy network on TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$, (50,100)(50,100)( 50 , 100 ), (100,50)(100,50)( 100 , 50 ), and (100,100)(100,100)( 100 , 100 ), and then directly apply the meta policy network to solve larger-sized problem instances that are never seen in the training stage (without any fine-tuning). The results are summarized in Table 5. The first block of Table 5 presents the results on U-TPP instances of sizes $\left(|M|,|K|\right)=(150,150)$, (200,200)(200,200)( 200 , 200 ), and (300,300)(300,300)( 300 , 300 ), and the second block presents the results on R-TPP instances with $|M|=150,|K|=150$ and $\lambda=0.99,0.95,0.9$.
It is demonstrated that our DRL-based approach still shows an advantage over the baseline heuristics when generalizing to U-TPP instances of sizes $\left(|M|,|K|\right)=(150,150)$ and (200,200)(200,200)( 200 , 200 ), while the largest instances for training are only of size $\left(|M|,|K|\right)=(100,100)$. However, we observe a drop in the zero-shot generalization performance on U-TPP instances of size over $\left(|M|,|K|\right)=(300,300)$ and R-TPP instances of size over $\left(|M|,|K|\right)=(150,150)$. In future work, we shall further explore the generalization and more efficient transfer learning approaches for larger-sized TPP instances.
@@ 8 Conclusion
[…]
This work was supported by the National Natural Science Foundation of China [Grants 61936009 and 72322021], and was supported by the Research and Development Project of Lenovo Group Ltd. and the Research and Development Project of CRSC Research & Design Institute Group Co., Ltd.

### v5

@@ 5.1 Policy Network Architecture
5.1.1 Input Embedding Module
Taking the bipartite graph representation of instance $U$ as input, the input embedding module exploits the raw features and topology of the bipartite graph to produce a high-dimensional embedding for each market node and product node. GNN is an effective framework that naturally operates on graph-structured data and learns relational information through message-passing between the nodes on the graph (Wu et al., 2021; Cappart et al., 2023). Therefore, we leverage GNNs in the input embedding module to embed the bipartite graph and capture the relations between markets and products.
Considering the feature heterogeneity of the bipartite graph representation, we first normalize the raw node and edge features by the product demands and linearly project them into the initial embeddings with a uniform dimension. Then, based on the structure of the input bipartite graph, we design a two-phase message-passing procedure that updates the initial node embeddings by aggregating information across the market nodes and product nodes: the first phase is performed to update the embeddings of product nodes, followed by the second phase that updates the embeddings of market nodes. Specifically, let $g_{k}^{\text{init}}$, $h_{m}^{\text{init}}$ and $e_{km}^{\text{init}}$ denote the initial embeddings of product node $k$, market node $m$, and edge $(k,m)$, respectively. The two-phase update is processed as follows:
@@ 5.1.1 Input Embedding Module
where $\operatorname{MLP}_{\mathcal{K}}$ and $\operatorname{MLP}_{\mathcal{M}}$ are multi-layer perceptrons (MLPs) for updating the product nodes and market nodes, respectively, and $\mathcal{N}\left(k\right)$ and $\mathcal{N}\left(m\right)$ denote the 1-hop neighbor node set of product node $k$ and market node $m$, respectively. Both phases can be viewed as a modification of the graph isomorphism network (GIN) (Xu et al., 2019) that further takes into account the bipartite structure and incorporates edge embeddings in the update. The updated product node embeddings $g_{k}^{0}$ for $k=1,2,\dots,|K|$ and market node embeddings $h_{m}^{0}$ for $m=0,1,\dots,|M|$ are then used in the market encoder and the decoder.
We remark that our design behind the two-phase embedding procedure is intuitive: in the first phase, we make each product node collect its supply information, including in which markets it is supplied, the coordinates of these markets, as well as the supply quantities and prices; and in the second phase, we make each market node collect higher-level information about its supplied products, including not only their total demands but also the information about how these products are supplied at other markets, which is extracted based on the product node embeddings from the first phase. Consequently, the product node embedding $g_{k}^{0}$ aggregates its global supply information, and the market node embedding $h_{m}^{0}$ collects more complex information such as the substitutive and complementary relations of product supply with other markets, which is important for the market selection decisions in route construction.
5.1.2 Market Encoder
It is empirically observed from prior works for routing problems that deeper information extraction for the city nodes, i.e., market nodes in our task, can potentially improve the capability of the policy network (Kwon et al., 2020; Joshi et al., 2022). Thus, we process the market node embeddings through a market encoder for further information extraction.
Following the Transformer (Vaswani et al., 2017) architecture, we process the market node embeddings through $N$ stacked attention layers. Each attention layer consists of two sublayers: an MHA layer that executes information aggregation for the market nodes and a node-wise MLP feed-forward layer to update the aggregated embeddings, where both sublayers add a residual connection (He et al., 2016) and batch normalization (BN) (Ioffe and Szegedy, 2015):
@@ 5.1.3 Decoder
The demand context $g_{\mathcal{K}}^{t}$ contains the information of remaining demand $d^{t}$. Third, the route context $o_{\pi}^{t}$ is the embedding for the partial route $\pi^{t}$, i.e., a sequence of market/depot nodes. We adopt an LSTM (Hochreiter and Schmidhuber, 1997) recurrent network, similar to the decoding RNN in Ptr-Net (Vinyals et al., 2015), to obtain the route context $o_{\pi}^{t}$:
@@ 5.2 Basic Training Algorithm
The policy network is updated based on the reward received after the solving process. As introduced in Section 3.2.4 and Section 3.2.5, given a TPP instance $U$, the policy network $\theta$ generates the final route $\pi$ with probability $p_{\theta}(\pi\mid U)=\prod_{t=1}^{T}p_{\theta}(a_{t}\mid s_{t})$ during training, where the associated traveling cost can be directly accessed, and the purchasing cost is obtained by solving a transportation problem. To align with the notations in DRL, we let $L(\pi\mid U)$ denote the loss of $\pi$, which is defined as the negative reward for the terminal state. We update the policy network $\theta$ to minimize the expectation of loss $L(\pi\mid U)$, using the REINFORCE algorithm (Williams, 1992) with baseline $b(U)$:
@@
where the baseline $b(U)$ can be seen as an estimation for the difficulty of instance $U$ to measure the relative advantage of route $\pi$ generated by the policy network. A well-defined baseline can significantly reduce the variance of gradients and accelerate training. We define the baseline $b(U)$ as the loss of the route obtained from a deterministic greedy rollout, i.e., where the action with the highest probability is always selected at each decision step, based on a baseline network $\theta^{\text{BL}}$ (Kool et al., 2019). Similar to the target Q-network in DQN (Mnih et al., 2015), the baseline network $\theta^{\text{BL}}$ is a copy of policy network $\theta$, but is fixed during each epoch to stabilize the baseline value. The baseline network $\theta^{\text{BL}}$ is periodically replaced by the latest policy network $\theta$ if the improvement of the policy is significant enough according to a paired t-test. Based on the gradient $\nabla_{\theta}\mathcal{L}(\theta\mid U)$, the parameters of policy network are updated using the Adam optimizer (Kingma and Ba, 2015). The basic training algorithm introduced above is outlined as Algorithm 1.
@@ 6 Meta-learning Strategy
Unfortunately, the meta-gradient involves a second-order derivative in $\frac{\partial\theta_{\text{in}}^{N}}{\partial\theta_{\text{in}}^{0}}$, which is significantly expensive in computation. Therefore, we instead employ a computationally efficient first-order approximation for the meta-gradient. The update for the meta policy network is approximated as moving towards the parameters of the final inner-loop policy network (Nichol et al., 2018):
@@
Algorithm 2 Meta-training for the policy network
@@ 7.1 Experimental Settings
7.1.1 Instances
We first randomly generate synthetic TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$, $(50,100)$, $(100,50)$, and $(100,100)$ for training and evaluation, where $|M|$ is the number of markets and $|K|$ is the number of products. The generation of U-TPP and R-TPP instances follows the standard procedure introduced in (Laporte et al., 2003), corresponding to Class 3 and Class 4 in the TPPLIB benchmark (Riera-Ledesma, 2012). For the U-TPP instance, $(|M|+1)$ integer coordinates (including the depot) are randomly generated within the $[0,1000]\times[0,1000]$ square according to a uniform distribution, and the traveling costs are defined as the truncated Euclidean distances. Each product $k\in K$ is supplied at $|M_{k}|$ randomly selected markets, where $|M_{k}|$ is uniformly generated in $\left[1,|M|\right]$. The price $p_{ik}$ for each product $k\in K$ and each market $i\in M_{k}$ is randomly sampled in $[1,10]$. The R-TPP instances are generated in the similar manner as U-TPP, with an additional limit $q_{ik}$ on the available supply quantities, which is randomly sampled in $[1,15]$ for each product $k\in K$ and each market $i\in M_{k}$. A parameter $\lambda$ is used to control the number of markets in a feasible solution through the product demand $d_{k}:=\left\lceil\lambda\max_{i\in M_{k}}q_{ik}+(1-\lambda)\sum_{i\in M_{k}}q_{ik}\right\rceil$ for $k\in K$, where $10<\lambda<$ 1. We employ $\lambda=\left\{0.99,0.95,0.9\right\}$ in our experiments, corresponding to the most difficult instances in the TPPLIB (Manerba et al., 2017).
@@ 7.1.1 Instances
7.1.2 Baseline Methods
We select several well-established TPP heuristics reported in (Manerba et al., 2017), including GSH, CAH, and TRH. A widely used and effective practice is to incorporate constructive heuristics (e.g., GSH and CAH) with TRH to remove redundant markets as soon as a solution is produced. Additionally, the solution can potentially be further improved by using a TSP solver to re-sequence the order of visited markets. In our experiment, we employ GSH and CAH, both followed by TRH and the TSP re-sequence, as the baselines, denoted as “GSH+TRH” and “CAH+TRH”, respectively.
Note that we do not include exact methods in the evaluation on synthetic instances since it can take several days to weeks to solve an evaluation instance set using exact methods (even for U-TPP). For a comparison of the optimality gap, we report the best-known solutions in the literature (Riera-Ledesma, 2012) on the TPPLIB benchmark instances to assess the optimality gaps of the baseline heuristics and our DRL-based approach.
@@ 7.1.2 Baseline Methods
Table 1: Training configuration.
7.1.3 Training and Inference of the Policy Network
@@
In the evaluation of our DRL-based approach, we execute the inference of the learned policy network with greedy decoding (similar to the greedy rollout baseline) and instance augmentation (Kwon et al., 2020). The produced route and purchasing plan can be directly used as an end-to-end solution to the TPP instance, denoted as “RL-E2E”. Similar to the baseline heuristics, this end-to-end solution can be further improved through a post-optimization procedure (TRH and TSP re-sequence), denoted as “RL+TRH”. For a fair comparison, the baseline heuristics and our approach are implemented on the same machine and environment.
@@ 7.2 Training Performance
The policy network is trained from scratch for U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$ and $(50,100)$. For brevity, we illustrate the training curve on U-TPP instances of size $\left(|M|,|K|\right)=(50,50)$ in Figure 4(a), which is plotted based on the average loss on the evaluation instance set. The stable convergence demonstrates that our policy network is capable of learning an effective policy for small-sized TPPs when trained from scratch using the basic training algorithm.
However, the policy network suffers serious training collapse when trained from scratch on large-sized U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(100,50)$ and $(100,100)$. As shown in Figure 4(b), the policy network trained from scratch on U-TPP instances of size $\left(|M|,|K|\right)=(100,50)$ is stuck in a high-loss region, and the loss completely fails to decrease throughout training. This is mainly because the state-action space is so large that the policy network cannot find an effective or even reasonable solution if the training starts from a random initialization. Therefore, we instead apply our meta-learning strategy to train the policy network for U-TPP and R-TPP instances with $\left(|M|,|K|\right)=(100,50)$ and $(100,100)$. As shown in Figure 4(b), the utilization of meta-learning strategy significantly promotes the training for an effective policy.
@@ 7.3 Results on Synthetic Instances
Following the training procedures described above, we train the policy network from scratch for U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$ and $(50,100)$, and apply the meta-learning strategy for training U-TPP and R-TPP instances of sizes $\left(|M|,|K|\right)=(100,50)$ and $(100,100)$. Then, the learned policy network is evaluated on the set of 1000 synthetic evaluation instances for each instance distribution. We report the metrics of 1) the average objective value and 2) the runtime (in seconds) for each evaluation instance set. The results are reported in Table 2 for U-TPP and Table 3 for R-TPP.
@@
Table 2: Results on synthetic U-TPP instances.
@@ 7.5 Generalization Performance on Large-Sized Instances
Finally, we examine the zero-shot generalization performance of our approach. We use the meta-learning strategy to train the policy network on TPP instances of sizes $\left(|M|,|K|\right)=(50,50)$, $(50,100)$, $(100,50)$, and $(100,100)$, and then directly apply the meta policy network to solve larger-sized problem instances that are never seen in the training stage (without any fine-tuning). The results are summarized in Table 5. The first block of Table 5 presents the results on U-TPP instances of sizes $\left(|M|,|K|\right)=(150,150)$, $(200,200)$, and $(300,300)$, and the second block presents the results on R-TPP instances with $|M|=150,|K|=150$ and $\lambda=0.99,0.95,0.9$.
It is demonstrated that our DRL-based approach still shows an advantage over the baseline heuristics when generalizing to U-TPP instances of sizes $\left(|M|,|K|\right)=(150,150)$ and $(200,200)$, while the largest instances for training are only of size $\left(|M|,|K|\right)=(100,100)$. However, we observe a drop in the zero-shot generalization performance on U-TPP instances of size over $\left(|M|,|K|\right)=(300,300)$ and R-TPP instances of size over $\left(|M|,|K|\right)=(150,150)$. In future work, we shall further explore the generalization and more efficient transfer learning approaches for larger-sized TPP instances.
@@
Table 5: Results on larger-sized instances.
@@ 8 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China [Grants 61936009 and 42327901], and was supported by the Research and Development Project of Lenovo Group Ltd. and the Research and Development Project of CRSC Research & Design Institute Group Co., Ltd.